Thinfilms morning report 3/14/05

james haydon jhaydon at
Mon Mar 14 09:16:16 PST 2005

Thinfilms Morning Report 3/14/2005



Gryphon:            ok

Innotec;            Shield change should be back up by late afternoon

STS Dep:            ok

Implanter:            noisy signal stil,l but ion gauge problem is fixed 
if you look at the end station ion gauge I will get to the noise problem 
next  week may take a little longer

Metalica:            ok

SCT:                ok

CPD:                ok                   

WbGen-CTB:            ok

WbGen-RHP:            ok

WbGen:-LHP:            ok

WbGaAs-RHP:            ok

WbGaSa-LHP:            ok

WbSilicide:            ok

WbDif: ok

WbNonmetal:            ok

WbNitride:            ok

WbMetal:            ok

Wbmiscres:            ok

WBOptres:            ok

WBSolvent:            ok

DI water:            ok

LN:                  ok


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <>

More information about the thinfilms mailing list